### v2

# testphase=new-or-1 MultiMat: Multimodal Program Synthesis for Procedural Materials using Large Multimodal Models

Abstract
Material node graphs are programs that generate the 2D channels of procedural materials, including geometry such as roughness and displacement maps, and reflectance such as albedo and conductivity maps. They are essential in computer graphics for representing the appearance of virtual 3D objects parametrically and at arbitrary resolution. In particular, their directed acyclic graph structure and intermediate states enable a modular, interpretable workflow for interactive appearance modeling. However, creating such graphs remains challenging and typically requires professional training. While recent neural program synthesis approaches attempt to simplify this process, they solely represent graphs as textual programs, failing to capture the inherently visual-spatial nature of node graphs that makes them accessible to humans. To address this gap, we present MultiMat, a multimodal program synthesis framework that leverages large multimodal models to process both visual and textual graph representations for improved generation of procedural material graphs. We train our models on a new dataset of production-quality procedural materials and combine them with a constrained tree search inference algorithm that ensures static correctness while efficiently navigating the program space. Our experimental results show that our multimodal program synthesis method is more efficient in both unconditional and conditional graph synthesis with higher visual quality and fidelity than text-only baselines, establishing new state-of-the-art performance.
1 Introduction
Procedural materials have become increasingly important in modern 3D content creation, offering artists greater control and flexibility in designing surface appearances for digital assets. Unlike traditional image-based textures, which are constrained by fixed resolutions and limited editability, procedural material modeling tools like Adobe Substance Designer (Adobe, 2025c) or Blender (Blender, 2025) leverage node-based graphs to generate textures programmatically. This enables resolution-independent execution, high-level parametric control, and non-destructive editing workflows that have proven valuable in industries such as game development, film production, and VR/AR applications (Musgrave et al., 2002). More specifically, a procedural material is defined as a directed graph where nodes represent texture generators (e.g., noise functions, patterns) or filtering operations (e.g., blurs, color adjustments), and edges encode the flow of data between these operations, ultimately producing the texture maps required by physically-based rendering (PBR) models (Pharr et al., 2016) (cf. Figure 1). However, the complexity of crafting these procedural material graphs presents a substantial barrier to entry, creating a pressing need for automated and semi-automated approaches to support material artists at all levels of proficiency.
With recent advances in neural program synthesis (Huynh & Lin, 2025), procedural material synthesis has become increasingly feasible. MatFormer pioneered this direction with a multi-stage transformer-based model for unconditional generation with Adobe Substance Designer (Guerrero et al., 2022). Building on this foundation, Hu et al. (2023) extended the approach to support conditional synthesis, enabling applications such as inverse rendering (Patow & Pueyo, 2003), i.e., generating procedural materials that match the appearance of captured or rendered images. More recently, VLMaterial demonstrated that large language models (Zhao et al., 2025) can effectively perform end-to-end procedural material synthesis (Li et al., 2025a). However, these approaches share a fundamental limitation: they generate node graphs as text-only programs without access to visual feedback during synthesis. This contrasts sharply with how human artists work, who create procedural materials by manipulating node graphs through an arguably more intuitive visual interface, as illustrated in Figure 1 (left). Without visual feedback, models must rely solely on textual representations to reason about complex spatial relationships and visual outcomes, a task that becomes increasingly difficult as material complexity grows. To address this limitation, we propose a novel multimodal program synthesis paradigm based on large multimodal models (Yin et al., 2024) that incorporates visual feedback throughout the generation process, more closely mirroring human creative workflows. We demonstrate that this approach, to which we refer as MultiMat, outperforms previous state-of-the-art methods (cf. §6). Our key contributions are as follows:
1.
We introduce MultiMat, a novel procedural material synthesis approach that incorporates visualizations of intermediate graphs, including node states, into its context. This multimodal feedback loop improves material quality substantially compared to text-only baselines.
2.
Investigating intermediate states enables real-time validation of each generated node. This allows us to develop a tree search algorithm that backtracks upon encountering invalid states, enabling more efficient inference than prior methods, which often produce invalid graphs.
3.
We implement a transpiler that converts between Adobe Substance Designer formats and a compact representation suitable for language modeling while supporting the complete feature set. This enables training on larger datasets and the generation of more complex materials than previous approaches, which examined only limited subsets of Designer’s capabilities.
2 Related Work
Large Language Models for Program Synthesis
Our work builds upon recent advances in neural program synthesis (Parisotto et al., 2017; Devlin et al., 2017; Thakoor et al., 2018; Ye et al., 2021; Ellis et al., 2021). Traditional program synthesizers require formal specifications and employ search or logical derivation to produce programs that provably satisfy these specifications (Alur et al., 2013). Recently, large language models have demonstrated impressive capabilities in this domain (Huynh & Lin, 2025; Li et al., 2025b; Lozhkov et al., 2024; Li et al., 2023b; Rozière et al., 2023; Fried et al., 2023; Li et al., 2022; Chen et al., 2021). However, current research predominantly targets high-resource programming languages such as Python, Java, or JavaScript (Zan et al., 2023; Huynh & Lin, 2025). In contrast, our work synthesizes graphics programs, which pose unique challenges due to domain-specific requirements and considerable data scarcity, establishing it as a distinct research area.
Graphics Program Synthesis
Deep learning approaches have shown strong performance in synthesizing graphics programs that compile to visual outputs (Ellis et al., 2018; 2019; Ganin et al., 2018). This progress has been accelerated by the emergence of large multimodal models, particularly vision-language models that bridge visual and textual domains (Alayrac et al., 2022; Liu et al., 2023; Belouadi et al., 2024b; Kulits et al., 2024; Li & Ellis, 2024; Kapur et al., 2025; Lin et al., 2025; Xu et al., 2025). The field encompasses both controlled experimental settings using domain-specific languages (Ellis et al., 2018; Tian et al., 2019; Sharma et al., 2018; Cámara et al., 2023; Kulits et al., 2024; Kapur et al., 2025; Wen et al., 2025) and practical applications. Notable examples include systems for generating scientific figures using TikZ (Belouadi et al., 2024a; b; 2025; Laurençon et al., 2024; 2024; Tong et al., 2024; Zhang et al., 2025) and automating data visualization (Mackinlay, 1986; Roth et al., 1994; Luo et al., 2021; Wu et al., 2024; Voigt et al., 2024). However, these approaches generate code designed for text-based editing and therefore do not face the unique circumstances of node graphs in procedural material synthesis that our work addresses.
Procedural Material Synthesis
Procedural material modeling is one of the most challenging domains in graphics program synthesis. The combination of lengthy, complex material programs and severe data scarcity creates unique obstacles for learning-based approaches (Li et al., 2025a; 2024). Existing methods primarily focus on inverse procedural material modeling by synthesizing graphs that reproduce a given target appearance (Hu et al., 2023) or unconditional generation to create diverse, novel materials without specific targets (Guerrero et al., 2022). A related line of work optimizes parameters of existing material graphs to match image targets by transpiling them into differentiable programs (Shi et al., 2020; Hu et al., 2022; Li et al., 2023a). As discussed in §1, previous generative approaches are limited to text-only representations, a limitation we address in this work.
3 Background on Procedural Materials
As indicated in §1, procedural materials are directed acyclic graphs , executed by a material engine to produce raster images representing the physical properties of materials. These so-called material maps define surface characteristics, e.g., albedo, roughness, or normal (tangent space orientation), that enable photorealistic rendering when applied to 3D objects, with their appearance controlled through a small set of high-level parameters (cf. Figure 1). The internal structure of a material graph comprises nodes connected by edges that define the flow of image data. Each node functions as either a generator that creates new image content or a filter that transforms existing images from upstream nodes. Common node operations include noise generation, blending, and mathematical transformations, which collectively produce intermediate image outputs . The behavior of each node is governed by parameters that may be discrete or continuous scalars or vectors, providing fine-grained control over the final material appearance.
Professional material authoring tools such as Blender and Adobe Substance Designer enable artists to construct and modify procedural material graphs through visual interfaces (cf. Figure 1). Users can interactively add or remove nodes and edges while adjusting node parameters to achieve desired visual effects. Among these tools, Adobe Substance Designer stands out for its particularly expressive node graph system, which MultiMat specifically targets. It offers advanced capabilities for creating complex material appearances through features like function graphs and pixel processors. Function graphs allow parameters to be controlled through custom operations on input values, while pixel processors enable users to define specialized computational graphs that operate on individual pixels using sequences of atomic mathematical operations. These sophisticated capabilities make automated procedural material synthesis a particularly challenging problem in this domain.
4 The MultiMat Model & Architecture
Figure 2 illustrates our complete model pipeline. At its core, MultiMat is a vision-language model, trained for synthesizing procedural material graphs. It accepts images as input for inverse procedural material synthesis and supports unconditional generation. Unlike previous approaches, MultiMat generates nodes topologically, ensuring each node precedes all nodes it connects to. This enables an iterative generation process detailed below that can provide continuous visual feedback to the model, verify the validity of intermediate outputs, and recover from errors automatically in certain cases.
4.1 Multimodal Program Synthesis
Given a partially generated material graph with nodes at generation step , the topological ordering of nodes allows for visualizing intermediate node states, similar to visual editing environments that target humans. This enables an iterative generation loop where MultiMat generates one node definition—including node parameters and connections to previous nodes—at a time that is processed accordingly before the generation continues. After generating node in an intermediate text format (cf. §5), we combine it with the existing node definitions and feed them to a transpiler, which compiles the intermediate representations back to a format the material engine understands. We then use the material engine to visualize the state of node . Upon successful transpilation and execution, is appended to the graph . This updated state, including the visualized intermediate outputs , is fed back to the model to generate the subsequent node (cf. Figure 2). If execution or transpilation fails, we discard the current and resample, or backtrack further in case of repeated errors (cf. §4.2). We explore two complementary approaches for representing and as multimodal programs to the model, as visualized in Figure 3:
- Mixed Conditioning
Starting with a textual representation of (cf. §5), we enhance each node with an additional field containing its visualized intermediate state. This creates a multimodal program where the model processes textual tokens interleaved with image patch embeddings (cf. Figure 3). To manage the increased context size from image embeddings, we omit node parameters (which are implicitly encoded in the visualizations) but explicitly include node output type information (e.g., grayscale or color) that the model cannot infer from the visualization alone.
- Graph Conditioning
This approach more closely mirrors human visual experience by conditioning MultiMat solely on a visualization of the entire graph with embedded intermediate visual outputs , as shown in Figure 3. The model generates subsequent node using only this complete visual context, without explicit access to underlying textual node definitions.
At their core, both approaches remain autoregressive language models, and MultiMat can be trained by minimizing a cross-entropy objective:
where is the expected token of in our intermediate text format at position , represents previous tokens, denotes the input conditions (which can be empty for unconditional generation), and represents the model parameters.
4.2 Incremental Tree Search
Another advantage of topological node ordering is the ability to validate node definitions incrementally during generation. By invoking our transpiler and material engine at each step, we can detect syntactic and semantic errors immediately rather than waiting until the entire graph is complete. When an erroneous node definition is encountered, we execute an adaptive backtracking strategy: first discarding and resampling the problematic node, and if errors persist, inferring deeper structural issues by reversing further back in the generation sequence. Specifically, we discard the most recently generated nodes, where represents the current backtracking iteration. This approach effectively transforms our generation process into an incremental tree search on a tree of valid and invalid nodes (cf. Figure 4), systematically exploring the solution space to discover valid programs. This incremental validation approach identifies invalid outputs much faster than previous approaches, which require sampling complete programs before validation can commence.
4.3 Automatic Error Repair
Through systematic analysis of failure cases, we identified recurring error patterns that could be repaired automatically: (1) removal of extraneous parameters that are specified for node types that do not support them, and (2) automatic insertion of conversion nodes to resolve type mismatches between connected nodes. For instance, when a color output is erroneously connected to a grayscale input, we automatically insert an appropriate grayscale conversion node. Conversely, when a grayscale output feeds into a color input, we insert a gradient map node to perform type conversion. These repair mechanisms increase the proportion of valid generations without requiring additional sampling steps.
5 Dataset
| Models | Size | Max Nodes | Feature Set | Program |
| MatFormer | Subset | Designer | ||
| Mat. (Cond) | Subset | Designer | ||
| VLMaterial | Limited | Blender | ||
| MultiMat | Complete | Designer |
1
Upper bound in complex filtering pipeline, actual could be less.
To support the training and evaluation of MultiMat, we collect procedural materials from Adobe’s Substance 3D Assets Repository (Adobe, 2025a). Unlike previous work that either focuses on basic graphs utilizing only a subset of Substance Designer features (e.g., lacking complex nodes such as pixel processors or function graphs; Guerrero et al., 2022; Hu et al., 2023) or targets other tools with more limited capabilities (Li et al., 2025a), our approach supports the complete feature set. This comprehensive coverage enables us to collect over 6 000 unique materials, substantially more than existing datasets. Table 1 summarizes key characteristics of our dataset compared to prior work.
Human-Readable Graph Representation
Substance Designer’s native file format (SBS) has not been designed for human readability, containing verbose XML structures, embedded binary data, legacy metadata, and other implementation details, which makes direct language modeling impractical. To address this, we develop a bidirectional transpiler that converts between SBS and a compact, human-readable YAML-based representation with topological node order, which we call CompactSBS. Unlike previous approaches that support only partial feature sets (Guerrero et al., 2022; Hu et al., 2023), our transpiler preserves the complete functionality of Substance graphs with programs that are, on average, over 80% shorter. Models operate exclusively in CompactSBS, with outputs transpiled back to SBS for execution. We provide representative examples in Figure 3 and complete program listings in Appendix A.
Graph Preprocessing
Our preprocessing pipeline standardizes graphs for the PBR workflow, focusing on five essential texture maps: base color, normal, roughness, metallic, and height. We trace backwards from these outputs to identify all contributing nodes, pruning unconnected components and other output maps. Graphs containing embedded bitmap graphics and SVGs are excluded to keep graphs fully procedural. We further filter out graphs exceeding 128 nodes and flatten hierarchical structures by inlining nested subgraphs and custom author dependencies into the main graph. Non-atomic nodes from the standard Substance Designer library remain as external references.
6 Experiments
We build MultiMat models upon the QWen2.5VL (7B) vision-language model which leverages a late fusion approach to combine image and text tokens (Bai et al., 2025). We train and evaluate separate models for unconditional generation (cf. §6.1) and inverse procedural material synthesis (cf. §6.2). We also conduct a human evaluation (cf. §6.3). Across all model variants, we maintain a consistent maximum sequence length of 8 192 tokens. The training setup consists of 5 epochs using AdamW (Loshchilov & Hutter, 2019), a learning rate of , and a batch size of 128. To ensure diversity in our generated outputs, we set the inference sampling parameters to a temperature of 0.8 and a top-p value of 0.95. We provide examples in Figure 6 and Appendix A. We ablate incremental tree search in §7.
6.1 Evaluation of Unconditional Generation
For unconditional generation, the mixed conditioning variant, MultiMat (Mixed), embeds node previews at resolution, resulting in 25 patch embeddings per image. For the graph conditioning variant, MultiMat (Graph), graph visualizations can utilize up to 6 144 tokens, with larger images downscaled to accommodate this limit. We generate 100 outputs per model for evaluation.
Baselines
For text-only procedural material synthesis, VLMaterial represents the current state-of-the-art approach. However, its Blender-specific training makes direct comparison with our method difficult. We therefore create VLMaterial (SBS) by retraining a VLMaterial-style model on our dataset for fair comparison. Unlike the objective in equation (1), VLMaterial is trained to generate complete graphs in a single pass. However, during inference, we can still validate nodes as they are generated and roll back upon detecting irreparable errors (cf. §4.2) or repair them after generation completes (cf. §4.3). This means the progression from VLMaterial (SBS) to MultiMat (Mixed) to MultiMat (Graph) represents a comparable, gradual shift from complete graph generation toward iterative node generation. Since VLMaterial (SBS) does not receive any images in the unconditional setting, we base it on the larger and more powerful text-only model QWen3 (8B; Yang et al., 2025a), giving it a slight advantage over our models. While graphics program synthesis research typically also benchmarks against proprietary large language models such as GPT-4o (OpenAI et al., 2024) or Claude 4 (Anthropic, 2025), which have demonstrated competitive performance in related domains (Belouadi et al., 2024a; b; 2025; Rodriguez et al., 2025), these models’ unfamiliarity with CompactSBS and inability to produce valid SBS output preclude their inclusion as baselines.
Metrics
Our multimodal task permits diverse evaluation schemes for automatic evaluation. To evaluate the visual quality of generated materials, we compute the Kernel Inception Distance (KID; Bińkowski et al., 2018), which compares the distribution of generated material maps with material maps from our dataset. To detect degenerate low KID scores due to memorization of training data (a legitimate concern given our relatively small dataset), we also calculate ROUGE-L scores (Lin, 2004) between the CompactSBS representation of our generated materials and the training set (with masked parameters). This metric computes the longest common subsequence and serves as an effective memorization indicator (Hans et al., 2024). Notably, we specifically require consecutive subsequences due to CompactSBS’s limited syntactic diversity, which could otherwise produce misleading matches. To measure efficiency, we introduce the Node Error Ratio (NER), defined as the average ratio between discarded nodes and the total number of generated nodes.
Results
| Models | KID ↓ | ROUGE-L ↓ | NER ↓ |
| VLMaterial (SBS) | |||
| MultiMat (Mixed) | |||
| MultiMat (Graph) |
| Models | DSim ↑ | CLIP ↑ | Style ↓ | KID ↓ | ROUGE-L ↓ | NER ↓ |
| VLMaterial (SBS) | ||||||
| MultiMat (Mixed) | ||||||
| MultiMat (Graph) | ||||||
| VLMaterial+ (SBS) | ||||||
| MultiMat+ (Mixed) | ||||||
| MultiMat+ (Graph) |
[…]
6.2 Evaluation of Conditional Generation
As in prior work (Hu et al., 2023; Li et al., 2025a), we train inverse MultiMat variants that learn to generate procedural materials from rendered images. These models follow the same training procedure as their unconditional counterparts, with one key modification: each training example is preceded by a rendering of itself, which adds 324 additional image patches to the model context. During inference, the model takes an image as input and generates a corresponding procedural material. We reserve 100 examples from our data as held-out test data for evaluation.
Baselines
Analogously to §6.1, we adapt VLMaterial for inverse rendering with SBS and use it as a baseline. Since an image input is now required for VLMaterial (SBS), we also base it on QWen2.5VL (7B) instead of QWen3 (8B) and train it using the same method as MultiMat.
Parameter Optimization
To further refine generated materials, we apply gradient-based optimization using differentiable rendering. This approach has proven effective for optimal parameter estimation (Shi et al., 2020; Hu et al., 2022; Li et al., 2023a; Hu et al., 2023). We employ DiffMat (Shi et al., 2020; Li et al., 2023a), a widely adopted differentiable renderer for Designer materials, to optimize the generated graphs against the input images. Models using this refinement step are denoted as MultiMat+ and VLMaterial,+ respectively.
Metrics
In addition to the metrics from §6.1, we evaluate reconstruction quality by rendering the generated materials and comparing them to the input images using perceptual similarity metrics. Specifically, we measure cosine similarity between CLIP image embeddings (Radford et al., 2021; Hessel et al., 2021), compute Style Loss loss (Style; Gatys et al., 2016) as the L1 distance between Gram matrices of VGG features, and calculate DreamSim (DSim; Fu et al., 2023), a learned perceptual similarity metric designed to align with human judgments.
Results
[…]
6.3 Human Evaluation
To corroborate our automatic evaluation results, we conduct a human evaluation. We employ comparative annotation (Thurstone, 1927) and focus on the image reconstruction/inverse rendering use case, which allows for intuitive human assessment (cf. Figure 6). Annotators receive triplets of rendered generated materials from VLMaterial+ (SBS), MultiMat (Mixed), and MultiMat (Graph) and identify which output best and least resembles the input image. Following Hu et al. (2023); Li et al. (2025a), we generate multiple programs () per model and image, selecting the result with the highest DreamSim score as the final candidate. We test 33 input materials from graphs filtered during preprocessing (e.g., due to excessive length), which represent particularly challenging cases. Eight expert annotators with extensive procedural material experience assess each triplet in randomized order. As shown in Figure 5, annotators rank VLMaterial+ (SBS) considerably lower than MultiMat+ (Mixed) and MultiMat+ (Graph), and prefer MultiMat+ (Graph) over MultiMat+ (Mixed). These findings align with our automatic evaluation rankings and demonstrate our approach’s effectiveness in generating perceptually similar materials.
7 Analysis & Discussion
| Input | VLMaterial+ (SBS) | MultiMat+ (Mixed) | MultiMat+ (Graph) |
Our comparisons demonstrate that model performance improves steadily as the degree of graph visualization increases, with MultiMat (Graph) achieving the highest performance overall (cf. Tables 2 & 3; Figure 5). This finding aligns with how humans interact with procedural materials—through visual node graph interfaces—and validates established UX design principles in this domain.
| Models | Deletion↓ | Conversion↓ |
| VLMaterial (SBS) | ||
| MultiMat (Mixed) | ||
| MultiMat (Graph) |
The qualitative examples in Figure 6 further illustrate this trend, with VLMaterial+ (SBS) struggling to generate faithful outputs, indicating that purely text-based approaches are not ideal for expressive node graph systems like Designer. This limitation persists even with more powerful base models, as our unconditional generation experiments confirm. Beyond architectural improvements, our tree search algorithm enables more efficient graph generation; without it, models may have to resort to sampling complete outputs for validation (the inference approach used by previous methods), which is expensive. For instance, disabling tree search causes NER of VLMaterial (SBS) to deteriorate further from 14.846 to 33.953, highlighting how our search strategy can improve inference without further training. The impact of automatic error repair is more nuanced, as shown in Table 4. Only approximately 1% of nodes generated by MultiMat contain hallucinated parameters, and fewer than 6.5% require conversion. In contrast, VLMaterial exhibits nearly double the scores for both repair mechanisms. This difference demonstrates that VLMaterial requires considerably more repair than our models and supports our claim that our models possess a better understanding of graph structures. Notably, since corrections are not fed back to the models, these results reflect their intrinsic generation capabilities.
8 Conclusion
We present MultiMat, a multimodal program synthesis framework and model suite that generates procedural materials by incorporating visual feedback throughout the generation process. Our key insight is that procedural material graphs are inherently visual-spatial programs, and treating them as such leads to substantial improvements over text-only approaches. By conditioning on visual intermediate states—either interleaved with text (mixed conditioning) or as complete graph visualizations (graph conditioning)—our models achieve consistent improvements over text-only baselines. Our incremental tree search algorithm further enhances generation efficiency by validating nodes as they are created and backtracking upon errors. While we demonstrate MultiMat specifically for procedural material synthesis, we hope its general principles will inspire further research at the intersection of computer graphics, program synthesis, and multimodal AI.
Future Work
The development of procedural material graph synthesis approaches is currently constrained by limited training data availability. We plan to address this challenge through self-learning techniques (He et al., 2020; Wei et al., 2021) that leverage our unconditional models to generate synthetic supervised training data by rendering outputs and subsequently training conditional models on this expanded data. Additionally, we aim to develop a unified model trained across multiple node graph systems to investigate potential transfer learning benefits (Pan & Yang, 2010). Beyond methodological advances, our models offer promising practical applications: conditional models could extract material graphs directly from photographic regions, while unconditional models could power intelligent auto-completion features in user interfaces. Furthermore, our methodology naturally extends to related domains such as vector graphics synthesis (Wu et al., 2023; Polaczek et al., 2025; Rodriguez et al., 2025; Yang et al., 2025b), where visual editing interfaces are similarly prevalent.
Limitations
Although our models and baselines use the same or similar base models, they generate graphs in fundamentally different ways, resulting in considerable differences in training efficiency. Text-only models like VLMaterial can process entire graphs as single training examples, whereas MultiMat must adapt the visual context for each individual node, effectively processing training examples one node at a time. This difference leads to much longer training times: while VLMaterial completes training in a few hours on 80GB GPUs, MultiMat models require several days on the same hardware despite being trained on a comparable number of tokens. Nevertheless, since the amount of procedural materials is very small (regardless of the dataset), training times remain within acceptable bounds in absolute terms, despite the relative differences between methods. Additionally, both approaches achieve a more similar throughput during inference.
Ethics Statement
We ensure that all procedural materials collected for model training are properly licensed and explicitly permit such usage, thereby preventing any copyright infringement. In adherence to this principle, we specifically exclude Substance 3D Community Assets (Adobe, 2025b) from our training data due to licensing restrictions. While we acknowledge the use of generative models in preparing this manuscript, their application is strictly limited to writing assistance, such as paraphrasing, spell checking, and synonym suggestions.
Acknowledgements
We thank the Adobe Substance 3D team for providing access to the Substance 3D Assets Repository and the Substance Automation Python API. We also thank our annotators for their valuable time. This work was conducted while the first author was an intern at Adobe Research, France.
Appendix A Additional Examples
In Figure 7 we provide additional qualitative examples. MultiMat+ (Mixed) consistently surpasses VLMaterial+ (SBS), while MultiMat+ (Graph) demonstrates the strongest results overall. Figure 8 complements Figures 6 & 7 by showcasing failure cases where our models struggle to produce faithful outputs, though notably, the outputs from MultiMat+ (Graph) and MultiMat+ (Mixed) still demonstrate superior representation of the input compared to VLMaterial (SBS). Beyond these conditional generation examples, Figure 9 presents unconditional samples generated by MultiMat (Graph), which exhibit high visual quality with realistic material properties. Adjacent to these rendered materials, we visualize their underlying material graphs in the same format used as model input. In Figure 10, we show a graph in CompactSBS representation to give an impression of the structure of our format.
| Input | VLMaterial+ (SBS) | MultiMat+ (Mixed) | MultiMat+ (Graph) |
| Input | VLMaterial+ (SBS) | MultiMat+ (Mixed) | MultiMat+ (Graph) |
| Render | Graph |